# A partition cover approach to tokenization

### v2

@@ Appendix A Tok’s relation to Unigram
[…]
$=\sum^{\mathbf{W}}_{W}\textsc{count}(W)\cdot(\log p(W^{*}_{i})+\dots+\log p(W^{*}_{|W^{*}|}))$ (notice $(\text{notice }partition(W)=|W^{*}|)$
[…]
@@
Iteration 2 random -1.431 random 1 0.0370 random random* = $\sum_{b\in W^{*}}p(b)$4.646
randose -1.431 randose 1 0.0370 randose random* = $\sum_{b\in W^{*}}p(b)$5.475
rosey -1.431 rosey 1 0.0370 rosey rosey* = $\sum_{b\in W^{*}}p(b)$3.743
randy -1.431 randy 1 0.0370 randy randy* = $\sum_{b\in W^{*}}p(b)$3.391
@@
[…]
@@
[…]
Table 6 shows the token pruning process of Unigram and Table 7 shows the greedy approach of GreedTok. For this scenario, GreedTok obtained a better solution with 10 partitions compared to Unigram’s 12. This scenario highlights the skew to select whole words when approaching tokenization from a pruning angle. If such scenarios are common, then we can expect GreedTok to obtain a better tokens per word ratio in our experiments (Section 5.1).
Appendix B Analyzing GreedTok’s Characteristics
After obtaining the token sets of GreedTok/BPE/Unigram from our experiments (Section 5), along three dimensions: 1) proportion of common tokens in a pairwise comparison, 2) proportion of whole words and 3) Unigram’s $\mathcal{L}$ objective (Appendix A).
@@ Appendix B Analyzing GreedTok’s Characteristics
To probe GreedTok’s token set, we analyze the common tokens from pairwise comparison of the tokenization methods investigated. From Fig. 5, we observe that BPE $\bigcap$ GreedTok share a large proportion of common tokens. The remaining difference can be observed in the greater proportion of Unigram $\bigcap$ GreedTok compared to Unigram $\bigcap$ BPE. The consistent results across the four corpora, and at different $|\mathbf{S}|=k$, implies that GreedTok contains the characteristics of BPE due to the high proportionality of common tokens. However, we require further investigation into whether GreedTok contains Unigram’s characteristics.
@@
As hypothesized in Appendix A, there are scenarios that will lead Unigram to select whole words, skipping the intermediate tokens that BPE may require. To further investigate, we approximate a 40,$0$ common word list using the most frequent words appearing in wiki, and then use it to compare to the token sets obtained. From Fig. 6, across the four corpora and at different $|\mathbf{S}|=k$, we observe that Unigram has the highest proportion of whole words, then followed by GreedTok and BPE respectively. With GreedTok selecting more whole words, compared to BPE, suggests that GreedTok may also exhibit the behaviour favored by Unigram.
@@
GreedTok (GT) $\mathcal{L}$ UN -4.15E08 -3.85E08 -3.69E08 -3.60E08 -3.53E08 PubMed -5.03E10 -4.71E10 -4.56E10 -4.47E10 -4.41E10
BPE $\mathcal{L}$4.37E08 -4.01E08 -3.83E08 -3.71E08 -3.63E08 -5.19E10 -4.83E10 -4.66E10 -4.56E10 -4.49E10
@@
Finally, another way that we can investigate the closeness of Unigram and GreedTok, is to examine the $\mathcal{L}$ of their token sets. From Table 8, we observe that GreedTok’s $\mathcal{L}$ is much closer to Unigram’s $\mathcal{L}$, relative to BPE’s.
@@
From the three analyses, there are two key findings. First, GreedTok’s compression ability can be explained by selecting a large proportion of tokens that BPE selects, and further improving it by selecting what BPE could not select, i.e. some tokens selected by Unigram. Second, there are indicators that GreedTok may display Unigram’s pre-training quality. This is observed from GreedTok having a higher proportion of common tokens and a closer $\mathcal{L}$ to Unigram relative to BPE. We investigate and confirm this in Section 5.2, where we conduct language pre-training on similar models trained on BPE versus GreedTok tokens.
Appendix C Additional Information
C.1 Additional Corpora Information
@@
UN [JBD17] is a collection of statements made by various member states during the General Debate on key geopolitical issues from 1946 to 2022. This corpus has a Creative Commons (CC) 0: Public Domain License.
@@
An extensive collection of English articles on a wide range of things, concepts, and people. We extract [Att15] the text from the database dump. [^7] Available at: https://dumps.wikimedia.org/. We also conduct a performance ablation with articles containing Chinese, Japanese and Korean (CJK) languages. The texts belonging to these articles are under CC BY-SA 4.0 and GNU Free Documentation Licenses.
@@
This corpus [PMH+23] is a filtered set of CommonCrawl [Com23] with stringent filtering and extensive deduplication. It was used to pretrain language models with 1B and 7B parameters. The corpus has a ODC-By 1.0 license.
@@
This corpus is built from DCLM [LFS+24] using a deduplication process that was used to build Zyda-2 [TGD+24]. This corpus is licensed under CC-by-4.
C.2 Additional Benchmark Information
We use the default settings of Language Model Evaluation Harness [GTA+24] for our benchmark evaluations.
@@ C.2 Additional Benchmark Information
A subset of easy questions from the Abstraction and Reasoning Corpus (ARC) dataset [CCE+18], comprising of 2,376 grade school level multiple-choice questions that were answered correctly by retrieval-based algorithm and a word co-occurrence algorithm. Each question has 4 answer options.
@@
Another subset of questions from the ARC dataset, the test set comprises 1,172 grade school level multiple-choice questions that were answered incorrectly by retrieval-based algorithm and a word co-occurrence algorithm. Similar to the easy subset, each question has 4 answer options.
@@
This challenge set [ZHB+19] evaluates the sentence completion ability in a multiple-choice manner. Given 4 possible answers, the correct answer completes the given sentence best. The test set contains 10,003 question sets.
@@
This dataset [MCKS18] contains multiple-choice questions that requires additional reasoning and knowledge in addition to the information given in the question and its 4 answer choices. We use the main set comprising of 500 test question sets.
@@
This dataset [BZB+20] contains binary-choice physical commonsense questions. Additional knowledge of how physical materials interact is required to successfully answer the questions. We use the validation set of 2,000 question sets.
@@
This dataset [JW17] contains multiple-choice science questions that were crowdscourced. Each question has 4 answer choices. We use the test dataset of 1000 question sets.
@@
This dataset [CLC+19] contains yes/no questions. Each question is accompanied by a corresponding text passage. We use the validation dataset of 3,270 triplets of question, passage, and answer.
@@
This dataset [RBG11] seeks to evaluate commonsense reasoning whereby given a premise and two plausible cause of effect, the correct answer is the option that is more plausible than the other. We use the validation set of 500 question sets.
@@
This dataset [PKL+16], abbreviated from LAnguage Modeling Broadened to Account for Discourse Aspects, contains text passages with the given task of predicting the last word of the target sentence. The task was crafted in a manner that the target word cannot be predicted by the target sentence, requiring information from other parts of the given text passage. We use the test dataset containing 5,153 text passages.
@@
This dataset [LXL+17] contains text passages with an associated multiple-choice question with four possible answer options. The questions were collected from English examinations in China. We use the high school level test set containing 3,498 passage-question sets.
Winogrande (Wino).
This dataset [SBBC19] follows the Winograd schema [LDM11], where the task is a fill-in-a-blank for a given sentence and two options, with additional commonsense reasoning required. We use the test dataset containing 1,767 question sets.
@@
We evaluate the bits/byte metric on Wikitext2 [MXBS16]. A lower value for this metric implies that less information (bits) is required to make a correct next token (byte) prediction.
Appendix D Additional Pseudocode
Previously, in our MIP (Section 4.2), a 1-based indexing system was used. However, for implementation convenience, we use a 0-based indexing system for our pseudocodes instead. Given an ordered sequence $\mathcal{S}$, such as array $A$, string $W$, and selected tokens $\mathbf{S}$, we use $\mathcal{S}_{i}$ to specify an element in the $i^{th}$ index of $\mathcal{S}$. However, for sequences, we use $\mathcal{S}_{i,j}$ to specify the elements from the $i^{th}$ index up to, but excluding, the $j^{th}$ of $\mathcal{S}$. For example, when $\mathcal{S}=\text{happy}$, we have $\mathcal{S}_{1}=\text{a}$ and $\mathcal{S}_{1,3}=\text{ap}$.
D.1 Computing $\mathcal{S}$ from $\mathcal{T}$
Given the Count function, corpus $\mathbf{W}$, candidate tokens $\mathbf{T}$, and an integer $k$, Algorithm 1 finds a set of tokens $\mathbf{S}$ that maximizes the objective function with the help of subroutines Algorithm 2 and Algorithm 3. The algorithm Algorithm 1 defines a couple of dictionaries $\mathbf{M}$, $\mathbf{P}$, $\mathbf{R}$, and $\mathbf{I}$ to track the problem state, then greedily picks the next best scoring token to cover words:
@@ D.1 Computing $\mathcal{S}$ from $\mathcal{T}$
    $\mathbf{M}$ maps each word $W$ to its state of cover, similar to the definition in Section 4.2
@@
    $\mathbf{P}$ maps each token $T$ to the set of its occurrences in the given word $W$, for all $W\in\mathbf{W}$, in a ($W$, $i$) pair, where $i$ is the position index of the start of the token occurrence
@@
    $\mathbf{I}$ maps each token $T$ to an index, which we use to update the state of cover for all $W\in\mathbf{W}$ at 20
@@
$\mathbf{M}(W)=(m^{W}_{0,1},\dots,m^{W}_{|W|-2,|W|-1})=(0,\ldots,0)=0^{|W|-1}\qquad\text{for all $W\in\mathbf{W}$}$
@@
4$\triangleright$ Positioning information of tokens in words
5:Initialize dictionary $\mathbf{R}:\mathbf{T}\to\mathbb{N}$ with $\mathbf{R}(T)=0$ for all $T\in\mathbf{T}$triangleright Token scores given current state
6:Initialize dictionary $\mathbf{I}:\mathbf{T}\to\mathbb{N}$ where $\mathbf{I}(T)=0$ for all $T\in\mathbf{T}$triangleright Token indices in $\mathbf{S}$
@@
8:Compute scores $\mathbf{R}(T)$ for each $T\in\mathbf{T}\setminus\mathbf{S}$ using Score on current state $\mathbf{M}$triangleright Algorithm 3
@@
10:Pop next best token $\tau=\operatorname*{argmax}_{T\in\mathbf{T}}$ from $\mathbf{Q}$ and add to back of $\mathbf{S}$
@@
13:     if new score $\hat{\mathbf{R}}(\hat{\tau})\neq\mathbf{R}(\hat{\tau})$ then $\triangleright$ Algorithm 3
14:         ${\mathbf{R}}(\hat{\tau})\leftarrow\hat{\mathbf{R}}(\hat{\tau})$
15:         Push $\hat{\tau},{\mathbf{R}}(\hat{\tau})$ back into $\mathbf{Q}$
@@
20:                  Update each entry of $\mathbf{M}(W)_{i,i+|\tau|-1}$ to $\mathbf{I}(\tau)$triangleright Update states of $\mathbf{M}(\tau)$
@@
2:Make a copy $\mathbf{M}^{\prime}$ of the state $\mathbf{M}$triangleright The original state remains unchanged
@@
6:         Add $\textsc{count}(W)\cdot|\{j\in\{i,\ldots,i+|T|\}:\mathbf{M}^{\prime}(W)_{j}=0\}|$ to $s$triangleright Only add score for non-zero entries
7:         Update each entry of $\mathbf{M}^{\prime}(W)_{i,i+|T|-1}$ to $1$triangleright Mark to avoid double counting; see Example 5
@@
The subroutine Algorithm 2 encapsulates a check of the validity of using a given token $T$ to cover $W$ at position $i$, primarily by observing if the non-start/end endpoint positions $i$ and $i+|T|$ were previously covered by some other token previously; if such a token is present, then $T$ cannot cover $W$ at position $i$. Meanwhile, the subroutine Algorithm 3 calculates the score contribution by token $T$, given the current state $\mathbf{M}$, while accounting for previous covers applied from chosen tokens in $\mathbf{S}$.

### v3

@@ Appendix D Tok’s relation to Unigram
$=\sum^{\mathbf{W}}_{W}\textsc{count}(W)\cdot\log W^{*}$ $(\text{group by words})$
$=\sum^{\mathbf{W}}_{W}\textsc{count}(W)\cdot\log\Pi^{W^{*}}_{W_{i}}p(W_{i})$ $(\text{choose best subword segmentation})$
$=\sum^{\mathbf{W}}_{W}\textsc{count}(W)\cdot(\log p(W^{*}_{i})+\dots+\log p(W^{*}_{|W^{*}|}))$ $(\text{notice }partition(W)=|W^{*}|)$
Notice that after grouping all similar words together, we get a weighted partition $(\log p(W_{i})+\dots+\log p(W_{|W^{*}|}))$. For Tok, it is $(1+\dots+1)=|W^{*}|$. Maximizing $\mathcal{L}$ is equivalent to minimizing negative log probability weighted partitions. Due to the presence of $\log$, $\max\mathcal{L}$ is equivalent to $\min\text{weighted }\textsc{Tok}$. We can see that Unigram favors frequently occurring subwords in a non-linear fashion. This relation is also noted in [50], where their proposed PathPiece tokenization algorithm optimizes for Tok from top-down pruning of BPE/Unigram shortlisted candidates.
Table 6: Top-down pruning solution for the given example using Unigram, removing tokens that results in the least decrease in $\mathcal{L}$, $\triangle\mathcal{L}$, in each iteration. The same results will also be obtained when optimizing for Tok from a top-down pruning approach. For iteration 2, $b$ denotes the singletons which are omitted. Final $\mathbf{S}=\{\text{``random''},\text{``randose''}\}$.
Iteration 1
random -1.505 random 1 0.0312 random random* $p(\text{``rand''})\cdot p(\text{``o''})\cdot p(\text{``m''}) -$2.056
@@
rosey -1.505 rosey 1 0.0312 rosey randy* $p(\text{`r'})\cdot p(\text{``ose''})\cdot p(\text{`y'}) -$1.806
randy -1.505 randy 1 0.0312 randy rosye* $p(\text{``rand''})\cdot p(\text{`y'}) -$0.727
@@
Iteration 2
random -1.431 random 1 0.0370 random random* = $\sum_{b\in W^{*}}p(b)$ -4.646
randose -1.431 randose 1 0.0370 randose random* = $\sum_{b\in W^{*}}p(b)$ -5.475
rosey -1.431 rosey 1 0.0370 rosey rosey* = $\sum_{b\in W^{*}}p(b)$ -3.743
randy -1.431 randy 1 0.0370 randy randy* = $\sum_{b\in W^{*}}p(b)$ -3.391
@@
Table 7: GreedTok’s solution for the given example, selecting tokens that results in the highest objective gain in each iteration. Final $\mathbf{S}=\{\text{``rand''},\text{``rosey''}\}$ or $\{\text{``rand''},\text{``ose''}\}$.
@@
We present a scenario where top-down pruning is sub-optimal for compression. Given $\mathbf{W}=\{\text{``random''},\text{``randose''},\text{``rosey''},\text{``randy''}\}$, each with a count of 1, and $\mathbf{T}=\mathbf{W}\bigcup\{\text{``rand''},\text{``ose''}\}$, we wish to select $\mathbf{S}\subset\mathbf{T}$, where $|\mathbf{S}|=k=2$ tokens.
Table˜6 shows the token pruning process of Unigram and Table˜7 shows the greedy approach of GreedTok. For this scenario, GreedTok obtained a better solution with 10 partitions compared to Unigram’s 12. This scenario highlights the skew to select whole words when approaching tokenization from a pruning angle. If such scenarios are common, then we can expect GreedTok to obtain a better tokens per word ratio in our experiments (Section˜5.1).
Appendix E Analyzing GreedTok’s Characteristics
After obtaining the token sets of GreedTok/BPE/Unigram from our experiments (Section˜5), along three dimensions: 1) proportion of common tokens in a pairwise comparison, 2) proportion of whole words and 3) Unigram’s $\mathcal{L}$ objective (Appendix˜D).
@@ Appendix E Analyzing GreedTok’s Characteristics
To probe GreedTok’s token set, we analyze the common tokens from pairwise comparison of the tokenization methods investigated. From Fig.˜5, we observe that BPE $\bigcap$ GreedTok share a large proportion of common tokens. The remaining difference can be observed in the greater proportion of Unigram $\bigcap$ GreedTok compared to Unigram $\bigcap$ BPE. The consistent results across the four corpora, and at different $|\mathbf{S}|=k$, implies that GreedTok contains the characteristics of BPE due to the high proportionality of common tokens. However, we require further investigation into whether GreedTok contains Unigram’s characteristics.
@@
As hypothesized in Appendix˜D, there are scenarios that will lead Unigram to select whole words, skipping the intermediate tokens that BPE may require. To further investigate, we approximate a 40,0000 common word list using the most frequent words appearing in wiki, and then use it to compare to the token sets obtained. From Fig.˜6, across the four corpora and at different $|\mathbf{S}|=k$, we observe that Unigram has the highest proportion of whole words, then followed by GreedTok and BPE respectively. With GreedTok selecting more whole words, compared to BPE, suggests that GreedTok may also exhibit the behaviour favored by Unigram.
@@
GreedTok (GT) $\mathcal{L}$
UN
-4.15E08 -3.85E08 -3.69E08 -3.60E08 -3.53E08
PubMed
-5.03E10 -4.71E10 -4.56E10 -4.47E10 -4.41E10
BPE $\mathcal{L}$ -4.37E08 -4.01E08 -3.83E08 -3.71E08 -3.63E08 -5.19E10 -4.83E10 -4.66E10 -4.56E10 -4.49E10
@@
Finally, another way that we can investigate the closeness of Unigram and GreedTok, is to examine the $\mathcal{L}$ of their token sets. From Table˜8, we observe that GreedTok’s $\mathcal{L}$ is much closer to Unigram’s $\mathcal{L}$, relative to BPE’s.
@@
From the three analyses, there are two key findings. First, GreedTok’s compression ability can be explained by selecting a large proportion of tokens that BPE selects, and further improving it by selecting what BPE could not select, i.e. some tokens selected by Unigram. Second, there are indicators that GreedTok may display Unigram’s pre-training quality. This is observed from GreedTok having a higher proportion of common tokens and a closer $\mathcal{L}$ to Unigram relative to BPE. We investigate and confirm this in Section˜5.2, where we conduct language pre-training on similar models trained on BPE versus GreedTok tokens.
Appendix F Additional Information
F.1 Additional Corpora Information
@@
UN [19] is a collection of statements made by various member states during the General Debate on key geopolitical issues from 1946 to 2022. This corpus has a Creative Commons (CC) 0: Public Domain License.
@@
An extensive collection of English articles on a wide range of things, concepts, and people. We extract [3] the text from the database dump. [^4] Available at: https://dumps.wikimedia.org/. We also conduct a performance ablation with articles containing Chinese, Japanese and Korean (CJK) languages. The texts belonging to these articles are under CC BY-SA 4.0 and GNU Free Documentation Licenses.
@@
This corpus [40] is a filtered set of CommonCrawl [10] with stringent filtering and extensive deduplication. It was used to pretrain language models with 1B and 7B parameters. The corpus has a ODC-By 1.0 license.
@@
This corpus is built from DCLM [29] using a deduplication process that was used to build Zyda-2 [52]. This corpus is licensed under $C$-by-4.
F.2 Model Pre-training Information
Our pre-training corpus is DCLM full-deduped dataset. For model training, we use the Dolomite Engine [33]. Our model architecture is a 40-layer Transformer [55], with embedding size 1536, MLP using SwiGLU activation [47] with intermediate size of 4096, and GQA [2] layers with 12 query heads and 4 pairs of KV-heads. We used a fixed context length of 4096 tokens and a batch size of $2^{22}\approx 4\mathrm{M}$ tokens. We train GreedTok and BPE tokenizers on approximately 20% of the dataset, randomly selected. For BPE, we use the popular implementation from Huggingface’s tokenizer API. [^6] Available at: https://github.com/huggingface/tokenizers. During training, we follow the same learning rate schedule as [51], training on the dataset in two phases: the Power Scheduler [51] in phase 1 and a learning rate with exponential decay in phase 2. However, our phases 1 and 2 use a similar data mixture, sampling the first 500M documents for phase 1, and the next 100M documents for phase 2, which always use 20% of the training iterations of phase 1. We run our experiments using NVIDIA H100 80GB HBM3 cluster, with 96 logical CPU count, training at a rate of $\sim$400B tokens/day.
Evaluating GTET is analogous to a setting where model training is compute-constrained, and users can use more text data for training. Conversely, evaluating GTEP is analogous to a setting where model training is data-constrained, and users have a limited amount of text for training. In the BPEM and GTET settings, we train the model for 125,000 and 25,000 iterations in phases 1 and 2 respectively. For GTEP, we take the model checkpoint of GTET at the 100,000th training iteration step, followed by an additional 20,000 training iterations in phase 2.
F.3 Additional Benchmark Information
We use the default settings of Language Model Evaluation Harness [16] for our benchmark evaluations.
@@ F.3 Additional Benchmark Information
A subset of easy questions from the Abstraction and Reasoning Corpus (ARC) dataset [7] contains 2,376 grade school level multiple-choice questions that were answered correctly with retrieval-based algorithm and a word co-occurrence algorithm. Each question has 4 answer options.
@@
Another subset of questions from the ARC dataset, the test set comprises 1,172 grade school level multiple-choice questions that were incorrectly answered with retrieval-based algorithm and a word co-occurrence algorithm. Similar to the easy subset, each question has 4 answer options.
@@
This challenge set [63] evaluates sentence completion in a multiple-choice setting. Given 4 possible answers, the correct answer completes the given sentence best. The test set contains 10,003 question sets.
@@
This dataset [32] contains multiple-choice questions that require additional reasoning and knowledge in addition to the information included in the question and its 4 answer choices. We use the main set containing 500 test question sets.
@@
This dataset [5] contains binary-choice physical commonsense questions. Additional knowledge of how physical materials interact is required to successfully answer the questions. We use the validation set of 2,000 question sets.
@@
This dataset [21] contains multiple-choice science questions that were crowdscourced. Each question has 4 answer choices. We use its test subset of 1000 question sets.
@@
This dataset [9] contains yes/no questions. Each question is accompanied by a corresponding text passage. We use the validation set of 3,270 triplets of question, passage, and answer.
@@
This dataset [43] seeks to evaluate commonsense reasoning, where given a premise and two plausible causes or effects, the correct answer is the option that is more plausible than the other. We use the validation set of 500 question sets.
@@
This dataset [38], abbreviated from LAnguage Modeling Broadened to Account for Discourse Aspects, contains text passages with the given task of predicting the last word of the target sentence. The task was crafted in a manner that the target word cannot be predicted by the target sentence, requiring information from other parts of the given text passage. We use its test set containing 5,153 text passages.
@@
This dataset [31] contains text passages with an associated multiple-choice question with four possible answer options. The questions were collected from English examinations in China. We use the high school level test set containing 3,498 passage-question sets.
Winogrande.
This dataset [45] follows the Winograd schema [28], where the task is a fill-in-a-blank for a given sentence and two options, with additional commonsense reasoning required. We use its test set containing 1,767 questions.
@@
We evaluate the bits/byte metric using Wikitext2 [34]. A lower value for this metric implies that less information (bits) is required to make a correct next-token (byte) prediction.
Appendix G Additional Pseudocode
Previously, in our MIP (Section˜4.1), a 1-based indexing system was used. However, for implementation convenience, we use a 0-based indexing system for our pseudocodes instead. Given an ordered sequence $\mathcal{S}$, such as array $A$, string $W$, and selected tokens $\mathbf{S}$, we use $\mathcal{S}_{i}$ to specify an element in the $i^{th}$ index of $\mathcal{S}$. However, for sequences, we use $\mathcal{S}_{i,j}$ to specify the elements from the $i^{th}$ index up to, but excluding, the $j^{th}$ of $\mathcal{S}$. For example, when $\mathcal{S}=\text{happy}$, we have $\mathcal{S}_{1}=\text{a}$ and $\mathcal{S}_{1,3}=\text{ap}$.
G.1 Computing $\mathcal{S}$ from $\mathcal{T}$
Given the Count function, corpus $\mathbf{W}$, candidate tokens $\mathbf{T}$, and an integer $k$, Algorithm˜1 finds a set of tokens $\mathbf{S}$ that maximizes the objective function with the help of subroutines Algorithm˜2 and Algorithm˜3. The algorithm Algorithm˜1 defines a couple of dictionaries $\mathbf{M}$, $\mathbf{P}$, $\mathbf{R}$, and $\mathbf{I}$ to track the problem state, then greedily picks the next best scoring token to cover words:
@@ G.1 Computing $\mathcal{S}$ from $\mathcal{T}$
    $\mathbf{M}$ maps each word $W$ to its state of cover, similar to the definition in Section˜4.1
@@
    $\mathbf{P}$ maps each token $T$ to the set of its occurrences in the given word $W$, for all $W\in\mathbf{W}$, in a ($W$, ii) pair, where $i$ is the position index of the start of the token occurrence
@@
    $\mathbf{I}$ maps each token $T$ to an index, which we use to update the state of cover for all $W\in\mathbf{W}$ at ˜20
@@
[…]
@@
[…]
5:Initialize dictionary $\mathbf{R}:\mathbf{T}\to\mathbb{N}$ with $\mathbf{R}(T)=0$ for all $T\in\mathbf{T}$ $\triangleright$ Token scores given current state
6:Initialize dictionary $\mathbf{I}:\mathbf{T}\to\mathbb{N}$ where $\mathbf{I}(T)=0$ for all $T\in\mathbf{T}$ $\triangleright$ Token indices in $\mathbf{S}$
@@
8:Compute scores $\mathbf{R}(T)$ for each $T\in\mathbf{T}\setminus\mathbf{S}$ using Score on current state $\mathbf{M}$ $\triangleright$ Algorithm˜3
@@
10:Pop next best token $\tau=\operatorname*{argmax}_{T\in\mathbf{T}}$ from $\mathbf{Q}$ and add to back of $\mathbf{S}$ $\triangleright$ Skip checking first token
@@
[…]
@@
20:      Update each entry of $\mathbf{M}(W)_{i,i+|\tau|-1}$ to $\mathbf{I}(\tau)$ $\triangleright$ Update states of $\mathbf{M}(\tau)$
@@
2:Make a copy $\mathbf{M}^{\prime}$ of the state $\mathbf{M}$ $\triangleright$ The original state remains unchanged
@@
6:   Add $\textsc{count}(W)\cdot|\{j\in\{i,\ldots,i+|T|\}:\mathbf{M}^{\prime}(W)_{j}=0\}|$ to $s$ $\triangleright$ Only add score for non-zero entries
7:   Update each entry of $\mathbf{M}^{\prime}(W)_{i,i+|T|-1}$ to 11 $\triangleright$ Mark to avoid double counting; see Example˜5
@@
The subroutine Algorithm˜2 encapsulates a check of the validity of using a given token $T$ to cover $W$ at position $i$, primarily by observing if the non-start/end endpoint positions $i$ and $i+|T|$ were previously covered by some other token previously; if such a token is present, then $T$ cannot cover $W$ at position $i$. Meanwhile, the subroutine Algorithm˜3 calculates the score contribution by token $T$, given the current state $\mathbf{M}$, while accounting for previous covers applied from chosen tokens in $\mathbf{S}$.